OpenStudy (anonymous):

Use the elimanation process to solve this equation x-2y=5 x+6y=13

OpenStudy (anonymous):

there on top of each other

OpenStudy (anonymous):

there is only one looking please help

terenzreignz (terenzreignz):

You know how the elimination process works?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Then where do you begin?

OpenStudy (anonymous):

im asking you this question'

terenzreignz (terenzreignz):

You did say you know how the elimination process works... Well, anyway, whenever you have an equality, say a = b h = k if you multiply both sides of an equality by a constant, equality holds: na = nb nh = nk Also, you can add the left-hand sides of the equalities, and it will be equal to the sum of the right-hand sides: a = b h = k a + h = b + k These are key to the elimination process.

terenzreignz (terenzreignz):

Take a look at your pair of equations: x - 2y = 5 x + 6y = 13 Of course, we can apply the second property so mentioned and right away add the left-hand sides and the right-hand sides of the equation, like so: (x - 2y) + (x + 6y) = 5 + 13 2x + 4y = 18 But that doesn't really help. The key is to use the first principle as well. In particular, multiplying (3) to the first equation before doing so, or multiplying (-1) to the second (or first) equation, before applying the second property would help a whole bunch... can you see why?